Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 566
Entry time: 18 Nov 2024, 09:41
|
Author: |
Frederik Wolff-Fabris |
Location: |
General |
Severity: |
Information |
Subject: |
Notes XO Integration Meeting 18.11.2024 |
|
|
General Announcement:
-
CTRL:
- Nothing to report.
DA:
- Nothing to report.
EEE:
- Nothing to report.
VAC:
- Nothing to report.
XPD:
- HAMP XGM at hard X-rays: emergency was triggered; after rebooting it worked normally.
DET:
- Nothing to report.
ITDM:
- Nothing to report.
XRO:
- Nothing to report.
LAS:
- Nothing to report.
- Low PP-laser pulse energy for FXE: not yet solved; needs investigation with AMPHOS - will talk to both SA1 instruments for use or time that intervention can be done.
UNSYS:
- Problem SA2 cell5: power supply distribution needed to be replaced last Thursday.
SPB:
- Nothing to report.
FXE:
- Nothing to report.
MID:
- Nothing to report.
HED:
- Not present.
SCS:
- Nothing to report.
SQS:
- Not present.
- Eric Boyd mention that the AIBS box at SQS will be replaced to a spare air box on Tuesday morning.
SXP:
- Nothing to report. |